OpenStudy (anonymous):

simplify. (7r^2-6t^2+2rt) + (-5rt+4t^2-2r^2)

OpenStudy (valpey):

\[(7r^2-6t^2+2rt) + (-5rt+4t^2-2r^2)\] \[7r^2-6t^2+2rt -5rt+4t^2-2r^2\] \[7r^2-2r^2+2rt -5rt+4t^2-6t^2\] \[(7-2)r^2+(2-5)rt+(4-6)t^2\] \[(5)r^2+(-3)rt+(-2)t^2\] \[5r^2-3rt-2t^2=(5r+At)(r+Bt)=5r^2+(A+5B)rt+(AB)t^2\]Can you solve for (A+5B)=-3 and AB = -2?

OpenStudy (anonymous):

No I'm not really sure how to do that ?

OpenStudy (valpey):

If A*B = -2 then A or B is either plus or minus 1 or plus or minus 2

OpenStudy (valpey):

One will be positive, the other negative.

OpenStudy (anonymous):

ok i got you so far

OpenStudy (valpey):

So which combination solves (A+5B)=-3?

OpenStudy (anonymous):

the -2?

OpenStudy (valpey):

What is -2?

OpenStudy (anonymous):

you said a * b = -2

OpenStudy (valpey):

Yes. What are A and B?

OpenStudy (anonymous):

are they -2?

OpenStudy (valpey):

-2*-2 = 4

OpenStudy (anonymous):

im not sure i don't like math

OpenStudy (anonymous):

u still there?